OpenStudy (anonymous):

Jacob walked for x miles and then cycled for y miles between his home and the library. the number of miles Jacob cycled was 3 miles more than twice the number of miles he walked. the total number of miles Jacob walked and cycled was 6 miles. how many miles did Jacob walk and cylcle?

OpenStudy (anonymous):

you will need to write equations using the given information. what i do is to first define any variables. usually, the variables are the quantities i want to know when i've finished the problem. can you define the variables and write equations using the given info?

OpenStudy (anonymous):

the variables would be x and y but I don't know how to write the equations

OpenStudy (anonymous):

x and y are lables. what do they represent? that's where you define them.

OpenStudy (anonymous):

x represents the miles he walked and y represents the miles he cycled

OpenStudy (anonymous):

okay, great! i ususally write statements to that effect, like: Let x = miles walked y = miles cycled that way it's clear to me and anyone else (like your teacher). Now re-read the problem and see if you can relate the info given using your variables.

OpenStudy (anonymous):

I can relate better to the problem but I still cant get the answer :(

OpenStudy (anonymous):

one step at a time... anty ideas on equations?

OpenStudy (anonymous):

not a clue... im really bad at math

OpenStudy (anonymous):

it's okay to say you don't have a clue but never say your bad at anything, especially something you want to learn. just say you're learning. what about this sentence... "the total number of miles Jacob walked and cycled was 6 miles." Can you make anything out of that?

OpenStudy (anonymous):

x+y=6?

OpenStudy (anonymous):

very good! now let's look at the sentence prior to that... "the number of miles Jacob cycled was 3 miles more than twice the number of miles he walked." can you get anything from this?

OpenStudy (anonymous):

y=x x 2 +3

OpenStudy (anonymous):

not sure what you're going for there. how about this: y = 2x+3

OpenStudy (anonymous):

okay now what do I do with that?

OpenStudy (anonymous):

so let's bring what we have together and think... we have the following: x + y = 6 and y = 2x + 6 any ideas on what we could do?

OpenStudy (anonymous):

sorry no..

OpenStudy (anonymous):

well, we could rewrite the first equation and instead of using y, we could use 2x + 6 instead, since they are equivalent. so we'd get: x + (2x + 6) = 6 can you solve this for x?

OpenStudy (anonymous):

x equals 0?

OpenStudy (anonymous):

sorry, it's supposed to be y = 2x + 3, not y = 2x + 6. so it would be x + (2x + 3) = 6. now solve for x.

OpenStudy (anonymous):

x equals 2

OpenStudy (anonymous):

I mean x=1

OpenStudy (anonymous):

there you go. no can you solvee for y? remember, y = 2x + 3

OpenStudy (anonymous):

y=5?

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

which graph is best for this answer?